A car with a mass of 1.2 x 103
kg is travelling at a speed of 12.4 m/s. What is the cars kinetic energy?
a) 92 J
b) 1.9 x 102
J
c) 7.7 x 103
J
d) 9.2 x 104
J
1
Expert's answer
2020-05-25T10:58:30-0400

The kinetic energy of a car

"KE=\frac{mv^2}{2}\\
=\frac{1.2\times 10^3\:\rm kg\times (12.4\:\ m/s)^2}{2}=9.2\times 10^4\: \rm J"

Answer: d)